Science and medicine. Martina leaves home at 9 A.M., bicycling at a rate of
24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina

Answer by ankor@dixie-net.com(12684)

Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours later,
John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina.
:
Let t = time that M is on the road
Then (t-2) = time J is on the road; (J leaves two hrs later)
:
When J catches up with M, they will have traveled the same distance.
Write a distance equation using this fact: (Distance = speed * time)
:
J's distance = M's distance
48(t-2) = 24t
48t - 96 = 24t
48t - 24t = +96
24t = 96
t = 96/24
t = 4 hrs after 9 A.M would be 1 P.M.
:
:
Check solution by seeing if they traveled the same distance:
24*4 = 96 mi
48*(4-2) = 96 mi
:
How about this? Did it make sense to you?
: